### v3

# Chem-FINESE: Validating Fine-Grained Few-shot Entity Extraction through Text Reconstruction

Abstract
Fine-grained few-shot entity extraction in the chemical domain faces two unique challenges. First, compared with entity extraction tasks in the general domain, sentences from chemical papers usually contain more entities. Moreover, entity extraction models usually have difficulty extracting entities of long-tailed types. In this paper, we propose Chem-FINESE, a novel sequence-to-sequence (seq2seq) based few-shot entity extraction approach, to address these two challenges. Our Chem-FINESE has two components: a seq2seq entity extractor to extract named entities from the input sentence and a seq2seq self-validation module to reconstruct the original input sentence from extracted entities. Inspired by the fact that a good entity extraction system needs to extract entities faithfully, our new self-validation module leverages entity extraction results to reconstruct the original input sentence. Besides, we design a new contrastive loss to reduce excessive copying during the extraction process. Finally, we release ChemNER+, a new fine-grained chemical entity extraction dataset that is annotated by domain experts with the ChemNER schema. Experiments in few-shot settings with both ChemNER+ and CHEMET datasets show that our newly proposed framework has contributed up to 8.26% and 6.84% absolute F1-score gains respectively [^1] The programs, data, and resources are publicly available for research purposes at: https://github.com/EagleW/Chem-FINESE..
Chem-FINESE: Validating Fine-Grained Few-shot Entity Extraction through Text Reconstruction
Qingyun Wang,  Zixuan Zhang,  Hongxiang Li,  Xuan Liu,   Jiawei Han,  Huimin Zhao,  Heng Ji University of Illinois at Urbana-Champaign {qingyun4,zixuan11,hanj,zhao5,hengji}@illinois.edu
1 Introduction
Millions of scientific papers are published annually [^2] https://esperr.github.io/pubmed-by-year/about.html, resulting in an information overload  Van Noorden (2014); Landhuis (2016). Due to such an explosion of research directions, it is impossible for scientists to fully explore the landscape due to the limited reading ability of humans. Therefore, information extraction, especially entity extraction of fine-grained scientific entity types, becomes a crucial step to automatically catch up with the newest research findings in the chemical domain.
Figure 1: Comparison of sentence reconstruction results from ground truth and InBoXBART Parmar et al. (2022). We highlight Complete Correct, Missed Entity, and Partially Correct Prediction with different color.
Despite such a pressing need, fine-grained entity extraction in the chemical domain presents three distinctive and non-trivial challenges. First, there are very few publicly available benchmarks with high-quality annotations on fine-grained chemical entity types. For example, ChemNER Wang et al. (2021a) developed the first fine-grained chemistry entity extraction dataset. However, their dataset is not released publicly. To address this issue, we collaborate with domain experts to annotate ChemNER+, a new chemical entity extraction dataset based on the ChemNER ontology. Besides, we construct another new fine-grained entity extraction dataset based on an existing entity typing dataset CHEMET Sun et al. (2021).
Figure 2: Type distributions for the training sets of ChemNER+ and CHEMET datasets. The Y-axis represents the number of mentions normalized by the mentions of the most frequent type. The X-axis represents the rank of types.
In addition, current entity extraction systems in few-shot settings face two main problems: missing mentions and incorrect long-tail predictions. One primary reason for missing mentions is that the sentences in scientific papers typically cover more entities than sentences in the general domain. For example, there are 3.1 entities per sentence in our ChemNER+ dataset, which is much higher than the 1.5 entities in the general domain dataset CONLL2003 Tjong Kim Sang and De Meulder (2003). As a result, it is more difficult for entity extraction models to cover all mentions in the input sentences. As shown in Figure 1, since the input has already included four chemical entities, InBoXBART model Parmar et al. (2022) completely misses the entity “room temperature”.
Furthermore, entity distributions in the chemical domain are highly imbalanced. As shown in Figure 2, we observe that the entity type distributions of ChemNER+ and CHEMET exhibit similar long-tail patterns. In few-shot settings, entities with long-tail types are extremely difficult to extract due to insufficient training examples. For example, as shown in Figure 1, InBoXBART mistakenly predicts the entity “aryl sulfamates” as catalyst, because its type has a frequency forty times lower than the predicted type (i.e., 4 vs 136). Moreover, the diverse representation nature of chemical entities—such as trade names, trivial names, and semi-systematic names (e.g., THF, iPrMgCl, 8-phenyl ring)—makes it even harder for models to generalize on these long-tail entities.
To address these challenges, we propose a novel Chemical FINe-grained Entity extraction with SElf-validation (Chem-FINESE). Specifically, our Chem-FINESE has two parts: a seq2seq entity extractor to extract named entities from the input sentence and a seq2seq self-validation module to reconstruct the original input sentence based on the extracted entities. First, we employ a seq2seq model to extract entities from the input sentence, since it does not require any task-specific component and explicit negative training examples Giorgi et al. (2022). We generate the entity extraction results as a concatenation of pairs, each consisting of an entity mention and its corresponding type, as shown in Figure 1.
One critical issue for seq2seq entity extraction is that the language model tends to miss important entities or excessively copy original input. For example, the seq2seq entity extraction results missed the type thermodynamic properties and generated “ligand screening” in Figure 1. However, the goal of information extraction is to provide factual information and knowledge comprehensively. In other words, if the model extracts knowledge precisely, readers should be able to faithfully reconstruct the original sentence using the extraction results. Inspired by such a goal, to evaluate whether the seq2seq entity extractor has faithfully extracted important information, we propose a novel seq2seq self-validation module to reconstruct the original sentences based on entity extraction results. As shown in Figure 1, the sentence reconstructed from the ground truth is closer to the original input than the sentence reconstructed from entity extraction results, which misses the reaction condition and introduces additional information that treated the “aryl sulfamates” as catalysts. Additionally, we introduce a new entity decoder contrastive loss to control the mention spans. We treat text spans containing entity mentions as hard negatives. For instance, given the ground truth entity “aryl sulfamates”, we will treat “aryl sulfamates at room temperature” as a hard negative.
Our extensive experiments demonstrate that our proposed framework significantly outperforms our baseline model by up to 8.26% and 6.84% absolute F1-score gains on ChemNER+ and CHEMET datasets respectively. Our analysis also shows that Chem-FINESE can effectively learn to select correct mentions and improve long-tail entity type performance. To evaluate the generalization ability of our proposed method, we also evaluate our framework on CrossNER Liu et al. (2021), which is based on Wikipedia. Our Chem-FINESE still outperforms other baselines in all five domains.
Our contributions are threefold:
  1. 1.
    We propose two few-shot chemical fine-grained entity extraction datasets, based on human-annotated ChemNER+ and CHEMET.
  2. 2.
    We propose a new framework to address the mention coverage and long-tailed entity type problems in chemical fine-grained entity extraction tasks through a novel self-validation module and a new entity extractor decoder contrastive objective. Our model does not require any external knowledge or domain adaptive pretraining.
  3. 3.
    Our extensive experiments on both chemical few-shot fine-grained datasets and the CrossNER dataset justify the superiority of our Chem-FINESE model.
2 Task Formulation
Following Giorgi et al. (2022), we formulate entity extraction as a sequence-to-sequence (seq2seq) generation task by taking a source document $\mathcal{S}$ as input. The model generates output $\mathcal{Y}$, a text consisting of a concatenation of $n$ fine-grained chemical entities $E_{1},E_{2},...,E_{n}$. Each mention $E_{i}$ includes the mention $\mu_{i}$ in the source document $\mathcal{S}$ and its entity type $\rho_{i}\in\mathcal{P}$, where $\mathcal{P}$ is a set containing all entity types. Specifically, we propose the following output linearization schema: given the input $\mathcal{S}$, the output is $\mathcal{Y}=\mu_{1}<\rho_{1}>,\mu_{2}<\rho_{2}>,...,\mu_{n}<\rho_{n}>$. We further illustrated this with an example:
$\mathcal{S}$: Through application of ligand screening, we describe the first examples of Pd-catalyzed Suzuki–Miyaura reactions using aryl sulfamates at room temperature.
$\mathcal{Y}$: ligand <Ligands>, Pd-catalyzed Suzuki–Miyaura reactions <Coupling reactions>, aryl sulfamates <Aromatic compounds>, room temperature <Thermodynamic properties>
Figure 3: Architecture overview. We use the example in Figure 1 as a walking-through example.
3 Method
3.1 Model Architecture
The overall framework is illustrated in Figure 3. Given the source document $S$, we first use a seq2seq model to extract fine-grained chemical entities. Then, we propose a new self-validation module to reconstruct the original input based on entity extraction results. Finally, we introduce a new entity decoder contrastive loss to reduce excessive copying. The entire model is trained with a combination of the supervised loss, the reconstruction loss, and the entity decoder contrastive loss.
3.2 Entity Extraction Module
Our entity extraction module follows a seq2seq setup Yan et al. (2021); Giorgi et al. (2022). Formally, we use the state-of-the-art coarse-grained chemical entity extractor InBoXBART Parmar et al. (2022) as the backbone. We model the conditional probability of extracting entities from source sequence $\mathcal{S}$ as
$p(\mathcal{Y}|\mathcal{S})=\prod_{t=1}^{T}p(y_{t}|\mathcal{S},y_{<t}),$ (1)
where the output $\mathcal{Y}$ has a length of $T$, and $y_{t}$ is the predicted token at time $t$ in the output $\mathcal{Y}$.
We supervise the entity extraction using the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{gen}}=\sum_{t=1}^{T}\log p(y_{t}|\mathcal{S},y_{<t}).$ (2)
3.3 Self-validation Module
Since a good information extraction system needs to extract entities faithfully, we propose a self-validation module to reconstruct the original sentence from the extracted entities to check whether the model overlooks any entities. Different from previous dual learning architectures Iovine et al. (2022), which use dual cycles or reinforcement learning to provide feedback, we use Gumbel-softmax (GS) estimator Jang et al. (2017) to avoid the non-differentiable issue in explicit decoding. Specifically, based on InBoXBART Parmar et al. (2022), we first pretrain a seq2seq self-validation module that takes in the entity extraction results $\mathcal{Y}$ and generates a reconstructed sentence $\mathcal{\hat{S}}$. We use our training set to pretrain the self-validation module. We fix the weight of the self-validation module after pretraining. In the training stage, the input embedding $\mathbf{H}_{t}$ of the self-validation module is given by:
$\mathbf{H}_{t}=\mathsf{GS}\left(p\left(y_{t}|\mathcal{S},y_{<t}\right)\right)\cdot\mathbf{E}_{v},$ (3)
where $\mathbf{E}_{v}$ is the vocabulary embedding matrix and $\mathsf{GS}$ is the Gumbel-softmax estimator. The total input embeddings for the self-reconstruction model is $\mathbf{H}=[\mathbf{H}_{1};\mathbf{H}_{2};...;\mathbf{H}_{T}]$.
The reconstruction loss is:
$\mathcal{L}_{\mathrm{recon}}=\sum_{\hat{t}=1}^{\hat{T}}\log p(\hat{s}_{\hat{t}}|\mathbf{H},\hat{s}_{<\hat{t}}),$ (4)
where the reconstructed sentence $\hat{\mathcal{S}}$ has a length of $\hat{T}$, and $\hat{s}_{\hat{t}}$ is the predicted token at time $\hat{t}$ in $\hat{\mathcal{S}}$.
3.4 Contrastive Entity Decoding Module
Entity extraction datasets in the scientific domain usually contain more entities for each sentence. From the initial experiments, we found that the entity extraction module tends to generate incorrect mentions by associating it with unrelated contexts to help the reconstruction of the self-validation module. For example, given the example in Figure 1, the baseline model generates “ligand screening” instead of “ligand”. Therefore, we introduce a new decoding contrastive loss inspired by Wang et al. (2023a) to suppress excessive copying. We construct negative samples by combining mentions with surrounding unrelated contexts. For example, we will consider “ligand screening, we describe the first examples” as a negative of entity “ligand”. We treat the original mention type pairs as the ground truth and maximize their probability with InfoNCE loss Oord et al. (2018):
$\begin{split}\mathcal{L}_{\mathrm{cl}}&=\frac{\exp{\left(x^{+}/\tau\right)}}{\sum_{i}\exp{\left(x^{-}_{i}/\tau\right)}+\exp{\left(x^{+}/\tau\right)}},\\ x^{+}&=\sigma(\mathrm{Avg}(\mathbf{W}_{x}\mathbf{\bar{H}}^{+}+\mathbf{b}_{x})),\\ x^{-}_{i}&=\sigma(\mathrm{Avg}(\mathbf{W}_{x}\mathbf{\bar{H}}^{-}_{i}+\mathbf{b}_{x})),\\ \end{split}$ (5)
where $\mathbf{\bar{H}}^{+}$ and $\mathbf{\bar{H}}_{i}^{-}$ are decoder hidden states from the positive and $i$-th negative samples, $\mathbf{W}_{x}$ is a learnable parameter, $\tau$ is the temperature, and $\mathrm{Avg}(*)$ denotes the average pooling function.
3.5 Training Objective
We jointly optimize the cross-entropy loss, reconstruction loss, and entity decoder contrastive loss:
$\mathcal{L}=\mathcal{L}_{\mathrm{gen}}+\alpha\mathcal{L}_{\mathrm{recon}}+\beta\mathcal{L}_{\mathrm{cl}},$ (6)
where $\alpha,\beta$ are hyperparameters that control the weights of the reconstruction loss and contrastive loss respectively.
Dataset Split #Pair
#$\mathbf{\overline{\#Token}}$
#$\mathbf{\overline{\#Entity}}$
Train
542
32.9
3.10
ChemNER+
Valid
100
39.9
4.57
Test
100
39.4
4.61
[…]
Train
6,561
37.8
1.57
CHEMET
Valid
520
31.6
2.15
Test
663
36.6
1.95
Table 1: Statistics of our dataset. #$\mathbf{\overline{\#Token}}$ denotes average number of words per sentence. #$\mathbf{\overline{\#Entity}}$ denotes average number of entities per sentence.
4 Benchmark Dataset
4.1 Dataset Creation
ChemNER+ Dataset.
Since the annotation of ChemNER dataset is not fully available online, we decide to create our own dataset, ChemNER+, based on available sentences from ChemNER Wang et al. (2021a) dataset. Following the schema of ChemNER, we ask two Chemistry Ph.D. students to annotate a new dataset, covering 59 fine-grained chemistry types with 742 sentences [^3] Human annotation details are in Appendix E..
CHEMET Dataset.
We construct a new fine-grained entity extraction dataset based on CHEMET Sun et al. (2021). For any entity in the training set that overlaps with the validation and testing sets, we replace its multi-labels with the most frequent types that appear in the validation and testing sets. For other entities, we replace the remaining types with their most frequent types that appeared in the training set. We merge the entity types with the same subcategory name in CHEMET Sun et al. (2021). The final dataset consists of 30 fine-grained organic chemical types.
Table 1 shows the detailed data statistics.
$k$-shot
6 9 12 15 18
RoBERTa
8.09
7.98
8.00
16.22
7.94
PubMedBERT
5.48
5.12
5.77
5.46
5.88
ScholarBERT
23.96
29.82
27.65
31.48
32.76
[…]
NNShot
0.99
1.43
2.39
1.61
2.45
StructShot
0.86
1.32
2.27
1.62
2.47
[…]
InBoXBART
26.23
27.89
28.83
33.64
30.39
+ Valid
32.40
31.13
33.64
35.31
36.44
+ Valid + CL
33.11 32.75 34.75 37.89 38.65
Table 2: micro-F1 (%) scores for ChemNER+ with few-shot settings. Valid is a model with a self-validation module. CL is a model with a decoder contrastive loss.
$k$-shot
6 9 12 15 18
RoBERTa
4.91
4.16
4.79
4.83
4.81
PubMedBERT
4.07
4.67
3.87
4.47
3.96
ScholarBERT
17.00
33.63
29.65
29.72
32.52
[…]
NNShot
4.23
4.03
4.14
5.27
4.76
StructShot
4.15
4.00
4.19
5.21
4.79
[…]
InBoXBART
29.93
29.57
31.76
36.16
37.52
+ Valid
32.74
34.09
33.30
40.81
38.37
+ Valid + CL
33.81 36.41 36.11
40.52
39.94
Table 3: micro-F1 (%) scores for CHEMET with few-shot settings.
4.2 Few-shot Setup
For each dataset, we randomly sample a subset based on the frequency of each type class. Specifically, given a dataset, we first set the number of maximum entity mentions $k$ for the most frequent entity type in the dataset. We then randomly sample other types and ensure that the distribution of each type remains the same as in the original dataset. We choose the values $6,9,12,15,18$ as the potential maximum entity mentions for $k$. The ChemNER+ and CHEMET few-shot datasets contain 52 and 28 types respectively.
5 Experiments
5.1 Baselines
We compare our model with (1) state-of-the-art pretrained encoder-based models including RoBERTa Liu et al. (2019) and models with domain adaptive training, such as PubMedBERT Gu et al. (2021) and ScholarBERT Hong et al. (2023). We then compare our model with the (2) few-shot baselines, including NNShot and StructShot Yang and Katiyar (2020) based on RoBERTa-base. Since we use InBoXBART Parmar et al. (2022) as our backbone, we also include (3) baselines for ablation. The hyperparameters, training and evaluation details are presented in Appendix A.
5.2 Overall Performance
Tables 2, 3 show that our models outperform baselines for few-shot settings by a large margin. Compared to the best pretrained encoder-based ScholarBERT, pretrained on 221B tokens of scientific documents, seq2seq models generally achieve higher performance in low-resource settings with fewer parameters, as shown in Table 11. We also observe that both NNshot and StructShot perform worse than their original baseline. At a closer look, we find that both methods miss many entities and mislabel unrelated phrases as entities. The primary reasons for this are twofold: first, the chemical domain’s entity mentions are more diverse and may only appear in the testing set; second, there are significantly more potential entity types than in traditional entity extraction tasks. Therefore, the two baselines cannot effectively utilize the nearest neighbor information and perform worse than our proposed methods. These results demonstrate that seq2seq models have a better generalization ability in few-shot settings.
$k$-shot
6 9 12 15 18
InBoXBART
36.96
38.22
38.34
47.91
42.84
+ Valid
45.07
45.28
41.56
48.15
46.15
+ Valid + CL
45.58
44.03
45.25 51.68 47.88
Table 4: Mention micro-F1 (%) scores for ChemNER+ with few-shot settings.
$k$-shot
6 9 12 15 18
InBoXBART
46.74
42.07
44.32
47.58
52.90
+ Valid
47.87
46.01
44.18
50.55
50.50
+ Valid + CL
48.96 49.83 47.03 50.61 54.10
Table 5: Mention micro-F1 (%) scores for CHEMET with few-shot settings.
Additionally, the self-validation variants significantly outperform the baseline InBoXBART, showing the benefit of the self-validation module in capturing mentions. Moreover, our self-validation module can effectively enhance the performance of the entity extraction module in extremely low-resource settings. In 6-shot scenarios for both ChemNER+ and CHEMET datasets, our model achieves impressive performance compared to ScholarBERT, which further verifies the effectiveness of the self-validation module. Finally, adding decoder contrastive loss helps the model perform significantly better in Table 2, suggesting that contrastive learning further helps the mention extraction quality by reducing excessive copying. Interestingly, we observe that decoder contrastive learning improves less in Table 3 than in Table 2, because the CHEMET contains fewer entities per sentence compared to the ChemNER+.
Figure 4: Average tokens in each mention for ChemNER+ and CHEMET datasets with few-shot settings.
Performance of Mention Extraction.
We calculate the mention F1 scores in Tables 4 and 5. In addition, we also test a fully unsupervised mention extraction based on AMR-Parser Fernandez Astudillo et al. (2020) [^4] Implementation details are in Appendix A.. The F1-scores are 38.22 and 45.33 for ChemNER+ and CHEMET, respectively. These results imply that the self-validation model generally improves the mention extraction accuracy. Moreover, adding decoder contrastive loss generally further bolsters the mention F1 score by reducing the number of tokens that appear in each mention, as shown in Figure 4.
$k$-shot
6 9 12 15 18
RoBERTa
2.04
2.05
2.05
0.00
2.05
PubMedBERT
2.05
0.00
0.00
2.13
0.00
ScholarBERT
0.00
9.28
4.71
0.00
6.90
[…]
InBoXBART
8.33
11.36
15.22
17.14
7.69
+ Valid
10.81
12.24
10.26
9.76
23.81
+ Valid + CL
26.19 23.91 23.26 19.05 25.00
Table 6: micro-F1 (%) scores for long-tail entity types ChemNER+ with few-shot settings.
$k$-shot
6 9 12 15 18
RoBERTa
0.00
0.00
0.00
0.00
0.00
PubMedBERT
0.00
0.00
0.00
0.00
0.00
ScholarBERT
0.00
0.00
0.00
0.00
0.00
[…]
InBoXBART
4.90
7.55
4.55
5.05
12.26
+ Valid
8.72 13.10
4.55
16.96
20.83
+ Valid + CL
7.07
11.32
8.33
5.15
23.01
Table 7: micro-F1 (%) scores for long-tail entity types CHEMET with few-shot settings. The encoder-based models fail to extract long-tail entity types for all few-shot settings. Compared to encoder-based models, seq2seq models can utilize label semantics in the generation procedure. Therefore, encoder-based models require more training data under few-shot settings.
Performance of Long-tail Entity.
To evaluate the performance of long-tail entities, we first rank entity types by their frequency. We then select the entity types that appear in the lower 50% and calculate the F1 scores of those types [^5] Entity frequency and selected types are in Appendix B.. The results are in Tables 6 and 7. Notably, our proposed methods greatly outperform the encoder-based baselines. Both the self-verification module and the decoder contrastive loss aid the entity extraction module in focusing on long-tail entities by creating a more balanced distribution of entity types. The major reason for the relatively low performance in Table 7 is that the differences between the types in CHEMET are not significant. The relatively stable performance of our model in Table 6 across increasing few-shot examples indicates that our model achieves satisfactory performance for long-tail entities, even with a limited training sample.
6 Analysis
6.1 Qualitative Analysis
Table 8 shows two typical examples from the 18-shot ChemNER+ dataset that illustrate how incorporating a self-validation module and decoder contrastive loss can improve the mention coverage and long-tail entity performance.
In the first example, the InBoXBART baseline fails to identify both “cyclophanes” and “polycycles”, probably because the input sentence contains too many entities. With the help of the self-validation module, the InBoXBART+Valid model successfully captures the first entity “cyclophanes”. However, it still cannot recognize “polycycles”. Additionally, both the baseline and the InBoXBART+Valid model mistakenly treat the entity “Suzuki cross-coupling and metathesis” and the entity “metathesis”, because those models excessively copy from the original sentence. In contrast, by adding the decoder contrastive loss, which uses the mentions with surrounding unrelated contexts as negatives, the model successfully separates the entity “Suzuki cross-coupling and metathesis” from the entity “metathesis”.
In the second example, both the baseline and the InBoXBART+Valid model predict a very long text span that treats three entities as a single entity. They also fail to capture “asymmetric catalysis” and “highly enantioselective process” as entities because their types have low frequency in the training set. With the help of decoder contrastive loss, the model reduces the excessive copying of the entity extraction module while trying to capture important entities as accurately as possible. Therefore, the model successfully classifies “asymmetric catalysis” as Catalysis correctly and also predicts “enantioselective process” as an entity.
InBoXBART
Several cyclophanes, polycycles, … have been synthesized by employing a combination of Suzuki cross-coupling and metathesis Coupling reactions.
[…]
+Valid
Several cyclophanes Heterocyclic compounds, polycycles, … have been synthesized by employing a combination of Suzuki cross-coupling and metathesis Organic reactions.
[…]
+Valid+CL
Several cyclophanes Heterocyclic compounds, polycycles Biomolecules, … have been synthesized by employing a combination of Suzuki cross-coupling Coupling reactions and metathesis Chemical properties.
[…]
Ground Truth
Several cyclophanes Aromatic compounds, polycycles Organic polymers, … have been synthesized by employing a combination of Suzuki cross-coupling Coupling reactions and metathesis Substitution reactions.
InBoXBART
… with the advantages of asymmetric catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Catalysis … the identification of a highly enantioselective process.
[…]
+Valid
… with the advantages of asymmetric catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Organometallic compounds … the identification of a highly enantioselective process
[…]
+Valid+CL
…with the advantages of asymmetric catalysis Catalysis (step and atom economy) in a rare example of an enantioselective cross coupling of a racemic electrophile bearing an oxygen leaving group Functional groups … the identification of a highly enantioselective process Chemical properties.
[…]
Ground Truth
… with the advantages of asymmetric catalysis Catalysis ( step and atom economy ) in a rare example of an enantioselective cross coupling Coupling reactions of a racemic electrophile Organic compounds bearing an oxygen leaving group Functional groups … the identification of a highly enantioselective process Catalysis.
Table 8: Examples showing how the self-validation module and entity decoder contrastive loss improves the model performance. We highlight Complete Correct, Missed Entity, and Partially Correct Prediction with different color. Compared to other baselines, our +Valid+CL successfully captures entities where other baselines miss.
6.2 Compatible with Other Few-shot Datasets?
CrossNER Dataset.
In the above experiments, we focus on the few-shot settings for chemical papers and prove the effectiveness of our proposed framework. To evaluate the generalization ability of our proposed framework on other domains, we conduct experiments on the CrossNER dataset Liu et al. (2021). The detailed statistics are in Table 9. We remove sentences without any entity. Because the CrossNER dataset is based on Wikipedia articles, we choose RoBERTa and ScholarBERT as encoder-based baselines. Additionally, we select BART-base Lewis et al. (2020) as the backbone for our ablation variations.
Dom. Train Valid Test #Type
#$\mathbf{\overline{\#Token}}$
#$\mathbf{\overline{\#Entity}}$
AI
100
350
430
14
31.5
4.42
Lit.
99
400
416
12
37.6
5.39
Mus.
100
380
465
13
41.4
7.05
Pol.
200
541
651
9
43.5
6.46
Sci.
200
450
543
17
35.8
5.62
Table 9: Statistics of CrossNER. Dom. denotes the domain of the dataset.
Model AI Lit. Mus. Pol. Sci.
RoBERTa
60.88
67.51
59.07
63.79
60.96
ScholarBERT
56.99
59.35
52.26
57.15
57.01
[…]
BART-base
59.20
66.90
62.78
67.99
62.18
+ Valid
61.84
67.97
60.94
67.22
62.40
+ Valid + CL
62.48 68.22 63.39 68.03 62.87
Table 10: F1 (%) scores for CrossNER.
Results.
As shown in Table 10, our model consistently produces the best F1 scores across all five domains of CrossNER without any external knowledge or domain adaptive pretraining. We observe that the model achieves the largest gain for the AI domain and the smallest gain for the politics domain. The major reason behind this is that AI domain contains the most informative entity types, which cover the key points of the sentence, including algorithm, task, etc. On the contrary, the politics domain contains many names of politicians and locations, which require background knowledge for the self-verification module to identify.
6.3 Remaining Challenges
Misleading Subwords.
We observe that the mention text can sometimes mislead the type predictions, especially if the type contains a subword from the mention. As a result, the model fails to identify the type correctly. For example, given the mention “unnatural amino acid derivatives”, our model focuses on the word “acid” and predicts the entity to be Organic acids instead of Organonitrogen compounds. The potential reason behind this is that the BART model incorrectly associates the “acid” in the mention with Organic acids. Such type errors might be incorporated into the decoder contrastive learning as additional hard negatives.
Fine-grained Type Classification.
The model tends to predict generic entity types instead of more fine-grained entity types. For instance, the model predicts the mention “Cs2CO3” as Inorganic compounds instead of Inorganic carbon compounds. This issue might come from annotation ambiguity in the training set. Additionally, the model predicts types that are not in the predefined ontology. For instance, the model labels “GK” as Genecyclic compounds instead of Enzymes. This error can possibly be solved by constraint decoding.
7 Related Work
Scientific Entity Extraction.
Entity extraction for scientific papers has been widely exploited in the biomedical domain Nguyen et al. (2022); Labrak et al. (2023); Cao et al. (2023); Li et al. (2023b); Hiebel et al. (2023) and the computer science domain Luan et al. (2018); Jain et al. (2020); Viswanathan et al. (2021); Shen et al. (2021); Ye et al. (2022); Jeong and Kim (2022); Hong et al. (2023). Despite this, fine-grained scientific entity extraction Wang et al. (2021a) in the chemical domain receives less attention due to the scarcity of benchmark resources. Most benchmarks in the chemical  Krallinger et al. (2015); Kim et al. (2015) only provide coarse-grained entity types. In this paper, we address this problem by releasing two new datasets for chemical fine-grained entity extraction based on the ChemNER schema Wang et al. (2021a) and CHEMET dataset Sun et al. (2021).
Few-shot Entity Extraction.
Few-shot learning attracts growing interest, especially for low-resource domains. Previous improvements for few-shot learning can be divided into several categories: domain-adaptive training by training the model in the same or similar domains Liu et al. (2021); Oh et al. (2022), prototype learning by learning entity type prototypes Ji et al. (2022); Oh et al. (2022); Ma et al. (2023), prompt-based methods Lee et al. (2022); Xu et al. (2023); Nookala et al. (2023); Yang et al. (2023); Chen et al. (2023b), data-augmentation Cai et al. (2023); Ghosh et al. (2023), code generation Li et al. (2023a), meta-learning de Lichy et al. (2021); Li et al. (2022); Ma et al. (2022), knowledge distillation Wang et al. (2021c); Chen et al. (2023a), contrastive learning Das et al. (2022), and external knowledge including label definitions Wang et al. (2021b), AMR graph Zhang et al. (2021), and background knowledge Lai et al. (2021). In contrast to these methods, our approach formulates the task in a text-to-text framework. In addition, we introduce a new simple but effective self-validation module, which achieves competitive performance without external knowledge or domain adaptive training.
Cycle Consistency.
Cycle consistency, namely structural duality, leverages the symmetric structure of tasks to facilitate the learning process. It has emerged as an effective way to deal with low-resource tasks in natural language processing. First introduced in machine translation He et al. (2016); Cheng et al. (2016); Lample et al. (2018); Mohiuddin and Joty (2019); Xu et al. (2020) to deal with the scarcity of parallel data, cycle consistency has been expanded to other natural language processing tasks, including semantic parsing Cao et al. (2019); Ye et al. (2019), natural language understanding Su et al. (2019); Tseng et al. (2020); Su et al. (2020), and data-to-text generation Dognin et al. (2020); Guo et al. (2020); Wang et al. (2023b). Recently, Iovine et al. (2022) successfully apply the cycle consistency to entity extraction by introducing an iterative two-stage cycle consistency training procedure. Despite these efforts, the non-differentiability of the intermediate text in the cycle remains unsolved, leading to the inability to propagate the loss through the cycle. To address this issue, Iovine et al. (2022) and Wang et al. (2023b) alternatively freeze one of the two models in two adjacent cycles. On the contrary, we introduce the gumbel-softmax estimator to avoid the non-differentiable issue. Additionally, we reduce the dual cycle training into end-to-end training to save time and computation resources.
8 Conclusion and Future Work
In this paper, we introduce a novel framework for chemical fine-grained entity extraction. Specifically, we target two unique challenges for few-shot fine-grained scientific entity extraction: mention coverage and long-tail entity extraction. We build a new self-validation module to automatically proofread the entity extraction results and a novel decoder contrastive loss to reduce excessive copying. Experimental results show that our proposed model achieves significant performance gains on two datasets: ChemNER+ and CHEMET. In the future, we plan to explore incorporating an external knowledge base to further improve the model’s performance. Specifically, we plan to inject type definition into the representation to facilitate the entity extraction procedure. We will also continue exploring the use of constraint decoding to further improve entity extraction quality.
9 Limitations
9.1 Limitations of Data Collections
Both ChemNER+ and CHEMET are based on papers about Suzuki Coupling reactions from PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/. Our fine-grained entity extraction datasets are biased towards the topics and ontology provided by ChemNER+ and CHEMET. For example, CHEMET only focuses on the organic compounds. The number of available sentences is limited by the original dataset and our annotation efforts. We currently only focus on the English sentences. We only test our model on chemical papers (i.e., ChemNER+ and CHEMET) and Wikipedia (CrossNER). In the future, we aim to adapt our model for categories in other languages.
9.2 Limitations of System Performance
Our few-shot learning framework currently requires defining the entity ontology and few-shot examples before performing any training and testing. Therefore, due to patterns in the pretraining set, our model might produce mention types that don’t align with our predefined ontology. For instance, it may generate Cyclopentadienyl compounds instead of the predefined type Cyclopentadienyl complexes. Furthermore, the pretrained model might emphasize language modeling over accurately identifying entire chemical phrases. For example, it might recognize Pd in the catalyst Pd(OAC)2 simply as a transition metal.
Acknowledgement
This work is supported by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897 , and by DOE Center for Advanced Bioenergy and Bioproducts Innovation U.S. Department of Energy, Office of Science, Office of Biological and Environmental Research under Award Number DESC0018420. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied of, the National Science Foundation, the U.S. Department of Energy, and the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Training and Evaluation Details
Avg. runtime # of Parameters
RoBERTa
16min
125M
PubMedBERT
18min
109M
ScholarBERT
19min
355M
InBoXBART
58min
139M
+Valid
56min
279M
+Valid+CL
59min
279M
Table 11: Runtimne (exclude CrossNER) and Number of Model Parameters
Our baselines and model are based on the Huggingface framework Wolf et al. (2020) [^7] https://github.com/huggingface/transformers. Our models are trained on a single NVIDIA A100 GPU. All hyperparameter settings are listed below. We optimize all models by AdamW Loshchilov and Hutter (2019). The runtime and number of parameters is listed in Table 11.
RoBERTa.
We train a RoBERTa-base model with 100 epochs and a batch size 32. The learning rate is $2\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. We use a linear scheduler for the optimizer.
PubMedBERT.
The PubMedBERT has the same model architecture as BERT-base with 12 transformer layers. The original checkpoint is pretrained on PubMed abstracts and full-text articles. We train a microsoft/BiomedNLP-PubMedBERT-base-uncased-abstract-fulltext model with 100 epochs and a batch size 32. The learning rate is $2\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. We use a linear scheduler for the optimizer.
ScholarBERT.
The ScholarBERT is based on the same architecture as BERT-large. The original checkpoint is pretrained on 5,496,055 articles from 178,928 journals. The pretraining corpus has 45.3% articles about biomedicine and life sciences. We train a globuslabs/ScholarBERT model with 100 epochs and a batch size 32. The learning rate is $2\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. We use a linear scheduler for the optimizer.
InBoXBART.
The InBoXBART is an instructional-tuning language model for 32 biomedical NLP tasks based on BART-base. We train the cogint/in-boxbart model with 100 epochs and a batch size 16. The learning rate is $10^{-5}$ with $\epsilon=1\times 10^{-6}$. During decoding, we use beam-search to generate results with a beam size 5. We use cosine annealing warm restarts schedule (Loshchilov and Hutter, 2017) for the optimizer.
InBoXBART+Valid.
We first pretrain the self-validation model, which is based on cogint/in-boxbart, on the training set. The learning rate for the self-validation module is $1\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. We use BLUE and ROUGE to select the best model. We then train the entity extraction model and the self-validation model jointly with cross-entropy $\mathcal{L}_{\mathrm{gen}}$ loss and reconstruction loss $\mathcal{L}_{\mathrm{recon}}$. The final loss is $\mathcal{L}=\mathcal{L}_{\mathrm{gen}}+5\cdot\mathcal{L}_{\mathrm{recon}}$. The learning rate is $5\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. During decoding, we use beam-search to generate results with a beam size 5. We use cosine annealing warm restarts schedule (Loshchilov and Hutter, 2017) for the optimizer.
InBoXBART+Valid+CL.
The final model is similar to InBoXBART+Valid. We retain the self-validation module and add a new decoder contrastive loss. The final loss is $\mathcal{L}=\mathcal{L}_{\mathrm{gen}}+0.2\cdot\mathcal{L}_{\mathrm{cl}}+5\cdot\mathcal{L}_{\mathrm{recon}}$. We randomly choose 5 negative samples for each instance. The learning rate is $5\times 10^{-5}$ with $\epsilon=1\times 10^{-6}$. During decoding, we use beam-search to generate results with a beam size 5. We use cosine annealing warm restarts schedule (Loshchilov and Hutter, 2017) for the optimizer.
AMR-based Mention Extraction.
We use AMR-parser Fernandez Astudillo et al. (2020) to extract mentions. We treat all text spans that are linkable to Wikipedia as mentions.
NNShot and StructShot.
We use the implementation from Ding et al. (2021) and choose RoBERTa-base as the language model.
Evaluation Metrics.
We use entity-level micro-F1 for all experiments. We use the library from nereval https://github.com/jantrienes/nereval.
Appendix B Dataset Details
We list the entity types of ChemNER+ and CHEMET below:
  • ChemNER+: Transition metals, Organic acids, Heterocyclic compounds, Organometallic compounds, Reagents for organic chemistry, Inorganic compounds, Thermodynamic properties, Aromatic compounds, Metal halides, Organic reactions, Alkylating agents, Organic compounds, Coupling reactions, Functional groups, Inorganic silicon compounds, Stereochemistry, Organohalides, Chemical properties, Catalysts, Free radicals, Alkaloids, Coordination chemistry, Ligands, Organophosphorus compounds, Reactive intermediates, Substitution reactions, Inorganic carbon compounds, Organonitrogen compounds, Biomolecules, Coordination compounds, Halogens, Chemical elements, Chlorides, Elimination reactions, Organic redox reactions, Inorganic phosphorus compounds, Organic polymers, Macrocycles, Cyclopentadienyl complexes, Substituents, Name reactions, Spiro compounds, Chemical kinetics, Organometallic chemistry, Catalysis, Organosulfur compounds, Ring forming reactions, Noble gases, Protecting groups, Addition reactions, Carbenes, Inorganic nitrogen compounds, Non-coordinating anions, Polymerization reactions, Carbon-carbon bond forming reactions, Isomerism, Enzymes, Oxoacids, Hydrogenation catalysts
  • CHEMET: Acyl Groups, Alkanes, Alkenes, Alkynes, Amides, Amines, Aryl Groups, Carbenes, Carboxylic Acids, Esters, Ethers, Heterocyclic Compounds, Ketones, Nitriles, Nitro Compounds, Organic Polymers, Organohalides, Organometallic Compounds, Other Aromatic Compounds, Other Hydrocarbons, Other Organic Acids, Other Organic Compounds, Other Organonitrogen Compounds, Other Organophosphorus Compounds, Phosphinic Acids And Derivatives, Phosphonic Acids, Phosphonic Acids And Derivatives, Polycyclic Organic Compounds, Sulfonic Acids, Thiols
The frequency for each type in the training data of both ChemNER+ and CHEMET are listed below:
  • ChemNER+: Organic compounds: 183, Coupling reactions: 171, Aromatic compounds: 136, Functional groups: 120, Heterocyclic compounds: 106, Catalysts: 70, Biomolecules: 66, Chemical elements: 64, Organohalides: 63, Transition metals: 56, Chemical properties: 55, Ligands: 55, Organic acids: 48, Thermodynamic properties: 43, Inorganic compounds: 43, Coordination compounds: 37, Stereochemistry: 33, Organometallic compounds: 33, Reagents for organic chemistry: 28, Coordination chemistry: 27, Organonitrogen compounds: 26, Organic reactions: 23, Organic polymers: 23, Substitution reactions: 21, Catalysis: 20, Organic redox reactions: 18, Reactive intermediates: 13, Substituents: 13, Halogens: 12, Addition reactions: 8, Chlorides: 6, Ring forming reactions: 6, Inorganic carbon compounds: 6, Enzymes: 6, Alkaloids: 4, Organophosphorus compounds: 4, Organosulfur compounds: 4, Oxoacids: 4, Elimination reactions: 3, Carbenes: 3, Inorganic phosphorus compounds: 2, Chemical kinetics: 2, Macrocycles: 2, Noble gases: 2, Organometallic chemistry: 2, Hydrogenation catalysts: 2, Metal halides: 1, Cyclopentadienyl complexes: 1, Inorganic nitrogen compounds: 1, Protecting groups: 1, Alkylating agents: 1, Polymerization reactions: 1
  • CHEMET: Other Organic Compounds: 1705, Ethers: 934, Other Aromatic Compounds: 882, Heterocyclic Compounds: 792, Alkanes: 528, Amides: 516, Other Organonitrogen Compounds: 501, Organometallic Compounds: 495, Esters: 440, Amines: 431, Ketones: 406, Polycyclic Organic Compounds: 375, Aryl Groups: 363, Organohalides: 312, Alkynes: 281, Alkenes: 266, Organic Polymers: 255, Other Hydrocarbons: 236, Other Organic Acids: 194, Other Organophosphorus Compounds: 97, Acyl Groups: 78, Nitriles: 77, Carboxylic Acids: 62, Sulfonic Acids: 37, Nitro Compounds: 26, Carbenes: 9, Phosphonic Acids And Derivatives: 4, Thiols: 2
We consider the following types as long-tail entity types for ChemNER+ and CHEMET. We list both the entity type and its frequency:
  • ChemNER+: Reactive intermediates: 13, Substituents: 13, Halogens: 12, Addition reactions: 8, Chlorides: 6, Ring forming reactions: 6, Inorganic carbon compounds: 6, Enzymes: 6, Alkaloids: 4, Organophosphorus compounds: 4, Organosulfur compounds: 4, Oxoacids: 4, Elimination reactions: 3, Carbenes: 3, Inorganic phosphorus compounds: 2, Chemical kinetics: 2, Macrocycles: 2, Noble gases: 2, Organometallic chemistry: 2, Hydrogenation catalysts: 2, Metal halides: 1, Cyclopentadienyl complexes: 1, Inorganic nitrogen compounds: 1, Protecting groups: 1, Alkylating agents: 1, Polymerization reactions: 1
  • CHEMET: Alkynes: 281, Alkenes: 266, Organic Polymers: 255, Other Hydrocarbons: 236, Other Organic Acids: 194, Other Organophosphorus Compounds: 97, Acyl Groups: 78, Nitriles: 77, Carboxylic Acids: 62, Sulfonic Acids: 37, Nitro Compounds: 26, Carbenes: 9, Phosphonic Acids And Derivatives: 4, Thiols: 2
Appendix C Evaluation on Whole Dataset
Model Precision Recall F1
In-BoXBART
55.73
43.28
48.72
+ Valid
57.49
45.77
50.97
+ Valid + CL
57.41
46.20 51.10
Table 12: micro-F1 for ChemNER+ with the whole training set.
We conduct fully supervised training on all training sets. The results are listed in Table 12 and 13. We observe that the self-validation module still improves the performance of the original InBoXBART for two datasets. We observe that the decoder contrastive loss further improves the model performance on ChemNER+. However, adding the entity decoder contrastive loss slightly decreases it. Because there are 6561 sentences in the CHEMET dataset, which is larger than the ChemNER+ dataset, the model with the self-validation module already performs very well. Additionally, since the CHEMET model contains fewer entities per sentence than the ChemNER+ dataset and these entities are all organic compounds separated away from each other, the entity decoder contrastive loss might introduce noise into the generation results, consequently decreasing the performance.
Model Precision Recall F1
In-BoXBART
64.94
41.62
50.73
+ Valid
70.09 42.16 52.65
+ Valid + CL
68.50
41.31
51.15
Table 13: micro-F1 for CHEMET with the whole training set.
Appendix D Scientific Artifacts
We list the licenses of the scientific artifacts used in this paper: PMC Open Access Subset Gamble (2017) [^8] https://www.ncbi.nlm.nih.gov/pmc/tools/openftlist/ (CC BY-NC, CC BY-NC-SA, CC BY-NC-ND licenses), Huggingface Transformers (Apache License 2.0), ChemNER (no license), CHEMET [^9] https://github.com/chenkaisun/MMLI1 (MIT license), RoBERTa (cc-by-4.0), PubMedBERT (MIT license), ScholarBERT (apache-2.0), BLEU [^10] https://github.com/cocodataset/cocoapi/blob/master/license.txt, ROUGE [^11] https://github.com/cocodataset/cocoapi/blob/master/license.txt, InBoXBART (MIT license), brat (MIT license), and nereval (MIT license). Our usage of existing artifacts is consistent with their intended use.
Appendix E Human Annotation
The instructions for human annotations can be found in the supplementary material. Human annotators are required to annotate the chemical compound entities mentioned either in natural language or chemical formulas and other chemical related terms including reactions, catalysts, etc. We recruit two senior Ph.D. students from the Chemistry department in our university to perform human annotations. We use brat Stenetorp et al. (2012) for all human annotations.
Appendix F Ethical Consideration
The Chem-FINESE model and corresponding models we have designed in this paper are limited to the chemical domain, and might not be applicable to other scenarios.
F.1 Usage Requirement
Our Chem-FINESE system provides investigative leads for few-shot fine-grained entity extraction for the chemical domain. Therefore, the final results are not meant to be used without any human review. However, domain experts might be able to use this tool as a research assistant in scientific discovery. In addition, our system does not perform fact-checking or incorporate any external knowledge, which remains as future work. Our model is trained on PubMed papers written in English, which might present language barriers for readers who have been historically underrepresented in the NLP/Chemical domain.
F.2 Data Collection
Our ChemNER+ sentences are based on papers from PMC Open Access Subset. Our annotation is approved by the IRB at our university. All annotators involved in the human evaluation are voluntary participants and receive a fair wage. Our dataset can only be used for non-commercial purposes based on PMC Open Access Terms of Use.